1. Solve the given differential equation by undetermined coefficients.
y'' − 8y' + 16y = 12x + 6

y(x) =_____

2. Solve the given boundary-value problem.
y'' + 7y = 7x, y(0) = 0, y(1) + y'(1) = 0

y(x) = _____
1
Expert's answer
2020-07-05T17:33:17-0400

1.The given equation can be written as "(D^2 \u2212 8D + 16)y = 12x + 6".

The auxiliary equation is, "m^2 \u2212 8m + 16 = 0." Solving we get, "m = 4 (twice)" . Thus, the complementary function is, "y_{c}=(c_{1}x+c_{2})e^{4x}". Let "y_{p} = Ax+B" be the trial particular solution. Then,


"y_{p}'' − 8y_{p}' + 16y_{p} = 12x + 6\\~\\
0-8A+16(Ax+B)=12x+6\\~\\
\therefore A= \frac{3}{4}, B= \frac{3}{4}\\~\\
\text{Hence,}~ y_{p}=\frac{3}{4}(x+1)"


Thus, the general solution is, "y=y_{c}+y_{p} = (c_{1}x+c_{2})e^{4x}+\frac{3}{4}(x+1)."


2.The given equation can be written as "(D^{2}+7)y=7x." The auxiliary equation is, "m^{2}+7=0." Solving we get, "m=\pm i\sqrt{7}". Thus, the complementary function is, "y_{c}=c_{1}\cos(\sqrt{7}x)+c_{2}\sin(\sqrt{7}x)." The particular solution is


"y_{p} = \dfrac{1}{D^{2}+7}7x\\~\\
~~~~~~~=\dfrac{1}{1+\frac{D^{2}}{7}}x\\~\\
~~~~~~~=\bigg(1+\dfrac{D^{2}}{7}\bigg)^{-1}x\\~\\ ~~~~~~~= x"


"y = c_{1}\cos(\sqrt{7}x)+c_{2}\sin(\sqrt{7}x)+x" is the general solution.

Given the boundary conditions, "y(0)=0; ~y(1)+y'(1)=0." Using these conditions we get,

"c_{1}=0\\~\\
c_{2}=\dfrac{-2}{\sqrt{7}\cos(\sqrt7)+\sin(\sqrt{7})}"


The general solution is,


"y = x-\dfrac{2\sin(\sqrt{7}x)}{\sqrt{7}\cos(\sqrt{7})+\sin(\sqrt{7})}"